# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 7.2 Concrete Atomic Functions
with proper symmetric $H(P)$ to be given in the theorems and corollaries below, satisfy (7.1) and hence are atomic functions for NEPv. Therefore, by Theorem 7.3, the NEPv Ansatz holds for them. These atomic functions are the same in form as the ones in subsection 4.2 but there are differences as detailed in Table 5. When inequality (4.1b) or (7.1b) become an equality, there is an important implication when it comes to verify the corresponding ansatz for the composition of atomic functions by a convex function $\phi$, namely, for equality (4.1b) or (7.1b), in general the corresponding partial derivative $\phi_{j}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{j}$ can be of any sign but is required to be nonnegative otherwise. We have seen this in Theorem 5.1 and will see it again in Theorem 8.1 later.
@@
$[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ $[\operatorname{tr}((P^{\operatorname{T}}AP)^{m})]^{s}$ NPDo $m=s=1$ $m\geq 2$ or $s>1$ $A\succeq 0$, (4.1b) an equality, (4.1b) an inequality, (4.1b) an inequality, $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$ or $\mathbb{O}^{n\times k}$. $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. $\mathbb{P}=\mathbb{O}^{n\times k}$. NEPv $m=s=1$ $m\geq 2$ or $s>1$ (7.1b) an inequality, (7.1b) an equality, (7.1b) an inequality, $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. $\mathbb{P}=\mathbb{O}^{n\times k}$. $A\succeq 0$, $\mathbb{P}=\mathbb{O}^{n\times k}$. * Integer $m\geq 1$, scalar $s\geq 1$ and $A$ is symmetric.
@@
for which $H(P)P-\mathscr{H}(P)\equiv P\big{[}m(D^{\operatorname{T}}P)^{m}\big{]}$ for $P\in\mathbb{O}^{n\times k}$. Then we have
$\operatorname{tr}(P^{\operatorname{T}}H(P)P)$ $=2m\,\operatorname{tr}((P^{\operatorname{T}}D)^{m})\quad\mbox{for$P\in\mathbb{O}^{n\times k}$};$ (7.7a)
@@
where $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}:=\{X\in\mathbb{O}^{n\times k}\,:\,X^{\operatorname{T}}D\succeq 0\}$. They, as argued in Remark 7.1, imply that (7.1) holds with $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=2(m-1)$, and thus $f(P)=\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ is an atomic function for NEPv. Furthermore, any solution $P$ to NEPv (6.3) with $H(P)$ in (7.6) such that $(P^{\operatorname{T}}D)^{m}$ is symmetric is a solution to the KKT condition (2.3) and vice versa.
@@
Equation (7.7a) can be straightforwardly verified. Now we prove (7.7b). Inequality (7.7b) for $m=1$ in fact holds for all $P\in\mathbb{O}^{n\times k}$. To see this, for $m=1$ and $P\in\mathbb{O}^{n\times k}$, since $\widetilde{P}^{\operatorname{T}}D\succeq 0$,
@@
by Lemma B.9, as expected. In general, suppose both $P^{\operatorname{T}}D\succeq 0$ and $\widetilde{P}^{\operatorname{T}}D\succeq 0$. Then
@@
Finally (7.7) implies (7.1), as a result of Remark 7.1. ∎
For any $s>1$, $[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ is well-defined for any $P\in\mathbb{R}^{n\times k}$ such that $\operatorname{tr}((P^{\operatorname{T}}D)^{m})\geq 0$. In particular, $[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ is well-defined for $P\in\mathbb{O}_{D+}^{n\times k}$. Combining Theorems 7.2 and 7.4, we obtain the following corollary.
@@
Let $D\in\mathbb{R}^{n\times k}$, integer $m\geq 1$, $s>1$, $g(P)=[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$, and $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. Let $H(P)$ be as in (7.6) and $H_{g}(P)=s\,[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s-1}H(P)$ for which $H_{g}(P)P-\partial g(P)/\partial P\equiv P\big{[}sm\,[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s-1}\,(D^{\operatorname{T}}P)^{m}\big{]}$. Then
@@
to get $\mathscr{H}(P)\equiv H(P)P$ for $P\in\mathbb{R}^{n\times k}$.
@@
They, as argued in Remark 7.1, imply that (7.1) holds with $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=2(m-1)$, and $\mathbb{P}=\mathbb{O}^{n\times k}$, and thus $f(P)=\operatorname{tr}((P^{\operatorname{T}}AP)^{m})$ is an atomic function for NEPv.
@@
They, as argued in Remark 7.1, imply that (7.1) holds with $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=2(sm-1)$, and $\mathbb{P}=\mathbb{O}^{n\times k}$, and thus $g(P)=[\operatorname{tr}((P^{\operatorname{T}}AP)^{m})]^{s}$ for $s>1$ is an atomic function for NEPv.
@@ 8 Convex Composition
We are interested in solving the same optimization problem on the Stiefel manifold $\mathbb{O}^{n\times k}$ as in (5.1) by Algorithm 1 and its accelerating variation in Algorithm 2 with convergence guarantee. In that regard, we stick to the initial setup at the beginning of section 5 up to the paragraph containing (5.3). We then go along a different path – the path of the NEPv approach. To that end, we will have to specify what $H(P)$, a symmetric matrix-valued function, to use for a given objective function $f=\phi\circ T$ in (5.1), assuming that a symmetric matrix-valued function has already been constructed for each component of $T(P)$.
@@ 8.1 All $P_{i}$ are the entire $P$
$\operatorname{tr}(P^{\operatorname{T}}H_{i}(P)P)$ $=\underaccent{\bar}{\gamma}_{i}\,f_{i}(P)\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (8.5a)
and given $\widehat{P}\in\mathbb{O}^{n\times k}$ and $P\in\mathbb{P}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
Given $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, suppose that (6.7) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\geq\operatorname{tr}(P^{\operatorname{T}}H(P)P)+\eta$. Write
@@
for which $H(P)P-\mathscr{H}(P)\equiv P\big{[}\mathscr{D}(P)^{\operatorname{T}}P\big{]}$ for $P\in\mathbb{O}^{n\times k}$, where, as in (5.14),
@@
Any solution $P$ to NEPv (6.3) with $H(P)$ in (8.6) such that $[\mathscr{D}(P)]^{\operatorname{T}}P$ is symmetric is a solution to the KKT condition (2.3) and vice versa. It can be seen that (8.5b) is an equality for $1\leq i\leq\ell$ and hence it does not need to require $\phi_{i}\geq 0$ for $1\leq i\leq\ell$. In addition to this, instead of treating each $H_{\ell+j}(P)$ separately, we can treat $H_{\ell+j}(P)$ for $1\leq j\leq t$ collectively all at once, as we did in (5.2), making all $\phi_{\ell+j}\geq 0$ unnecessary as well. A much more improved version of Theorem 8.1 is stated as Theorem 8.2 below, according to which, the best $Q_{i}$ at Line 4 of Algorithm 1 when applied to $\phi\circ T_{1a}$ is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}\mathscr{D}(P^{(i)})$. A special case of $T_{1a}$ is: $t=0$, $k=1$, $P_{i}=\boldsymbol{p}$ (a unit vector) for $1\leq i\leq\ell$, which gives the main problem of [5] (in the paper, $\phi(\boldsymbol{x})=\sum_{i=1}^{\ell}\psi_{i}(x_{i})$ for $\boldsymbol{x}=[x_{1},x_{2},\ldots,x_{\ell}]^{\operatorname{T}}$ with each $\psi_{i}$ being a convex function of a single-variable).
@@
Consider $f=\phi\circ T_{1a}$, and let $\mathscr{D}(P)$ as in (5.14) and $H(P)$ be as in (8.6). Given $\widehat{P}\in\mathbb{O}^{n\times k},\,P\in\mathbb{O}^{n\times k}$, let $\widetilde{P}=\widehat{P}Q$ where $Q$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{D}(P)$. If
@@
$T_{2a}\,:\,P\in\mathbb{O}^{n\times k}\to T_{2a}(P):=\begin{bmatrix}\|P^{\operatorname{T}}A_{1}P\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P^{\operatorname{T}}A_{\ell}P\|_{\operatorname{F}}^{2}\\ \|P^{\operatorname{T}}D_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P^{\operatorname{T}}D_{t}\|_{\operatorname{F}}^{2}\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (8.8)
@@
for which $\mathscr{H}(P)\equiv H(P)P$ for $P\in\mathbb{R}^{n\times k}$. If all $A_{i}\succeq 0$, then, by Theorem 7.5, the consistency conditions in (8.5) are satisfied with $\mathbb{P}=\mathbb{O}^{n\times k}$, $Q=I_{k}$, $\underaccent{\bar}{\alpha}=2$, $\underaccent{\bar}{\beta}_{i}=2$ for $1\leq i\leq\ell$ and $\underaccent{\bar}{\beta}_{\ell+j}=0$ for $1\leq j\leq t$. Note, for the example, (8.5b) for $\ell+1\leq i\leq\ell+t$ are equalities and hence Theorem 8.1 requires $\phi_{i}\geq 0$ for $1\leq i\leq\ell$ only. Optimization problem (5.1) with $T=T_{2a}$ for $t=0$ and $\phi(\boldsymbol{x})=\sum_{i=1}^{\ell}x_{i}$ gives the key optimization problem in the uniform multidimensional scaling (UMDS) method [79].
Comparing the conclusion here and that of Example 5.2 (for all $P_{i}=P$), we don’t require $\phi_{\ell+j}\geq 0$ for $1\leq j\leq t$ here, everything else being the same. In particular, both require $A_{i}\succeq 0$ for all $i$. Next, we explain how to make the NEPv approach work on $f=\phi\circ T_{2a}$ even if some $A_{i}\not\succeq 0$, yielding yet another example for which the NEPv approach works but the NPDo approach may not. Let $\delta_{i}\in\mathbb{R}$ such that $\widehat{A}_{i}=A_{i}-\delta_{i}I\succeq 0$ for $1\leq i\leq\ell$. This is always possible by letting $\delta_{i}$ be some lower bound of the eigenvalues of $A_{i}$, and numerically, $\delta_{i}$ can be estimated cheaply [81]. Notice that
$\operatorname{tr}((P^{\operatorname{T}}A_{i}P)^{2})=\operatorname{tr}((P^{\operatorname{T}}\widehat{A}_{i}P)^{2})+2\delta_{i}\operatorname{tr}(P^{\operatorname{T}}\widehat{A}_{i}P)+k\delta_{i}^{2},$
@@
$\widehat{T}_{2a}\,:\,P\in\mathbb{O}^{n\times k}\to\widehat{T}_{2a}(P):=\begin{bmatrix}\operatorname{tr}((P^{\operatorname{T}}\widehat{A}_{1}P)^{2})\\ \vdots\\ \operatorname{tr}((P^{\operatorname{T}}\widehat{A}_{\ell}P)^{2})\\ \operatorname{tr}(P^{\operatorname{T}}\widehat{A}_{1}P)\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}\widehat{A}_{\ell}P)\\ \|P^{\operatorname{T}}D_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P^{\operatorname{T}}D_{t}\|_{\operatorname{F}}^{2}\\ \end{bmatrix}\in\mathbb{R}^{2\ell+t},$ (8.10)
@@
$T_{3a}\,:\,P\in\mathbb{O}^{n\times k}\to T_{3a}(P):=\begin{bmatrix}\operatorname{tr}((P^{\operatorname{T}}A_{1}P)^{m_{1}})\\ \vdots\\ \operatorname{tr}((P^{\operatorname{T}}A_{\ell}P)^{m_{\ell}})\\ \operatorname{tr}((P^{\operatorname{T}}D)^{m_{\ell+1}})\end{bmatrix}\in\mathbb{R}^{\ell+1},$ (8.11)
@@
for which $H(P)P-\mathscr{H}(P)\equiv P\big{[}\phi_{\ell+1}(T_{3a}(P))\,m_{\ell+1}\,(D^{\operatorname{T}}P)^{m_{\ell+1}}\big{]}$ for $P\in\mathbb{O}^{n\times k}$. Any solution $P$ to NEPv (6.3) with $H(P)$ in (8.12) such that $(D^{\operatorname{T}}P)^{m_{\ell+1}}$ is symmetric is a solution to the KKT condition (2.3). Suppose all $A_{i}\succeq 0$ and let $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. Then the consistency conditions in (8.5) are satisfied with $\underaccent{\bar}{\alpha}=2$, $\underaccent{\bar}{\beta}_{i}=2(m_{i}-1)$ for $11\leq i\leq\ell+$ 1, by Theorems 7.4 and 7.5. In applying Theorem 8.1 with $T_{0}=T_{3a}$ we need $\phi_{i}\geq 0$ for $11\leq i\leq\ell+$ 1 because now we are not sure if any of the inequalities in (8.5b) is an equality. Lastly, the best $Q_{i}$ at Line 4 of Algorithm 1 when applied to $\phi\circ T_{3a}$ is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}D$ and vice versa.

### v4

@@ 7.2 Concrete Atomic Functions
with proper symmetric $H(P)$ to be given in the theorems and corollaries below, satisfy (7.1) and hence are atomic functions for NEPv. Therefore, by Theorem 7.3, the NEPv Ansatz holds for them. These atomic functions are the same in form as the ones in subsection 4.2 but there are differences as detailed in Table 5. When inequality (4.1b) or (7.1b) become an equality, there is an important implication when it comes to verify the corresponding ansatz for the composition of atomic functions by a convex function $\phi$, namely, for equality (4.1b) or (7.1b), the corresponding partial derivative $\phi_{j}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{j}$ can be of any sign but in general is required to be nonnegative otherwise. We have seen this in Theorem 5.1 and will see it again in Theorem 8.1 later.
@@
$[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ $[\operatorname{tr}((P^{\operatorname{T}}AP)^{m})]^{s}$ NPDo $m=s=1$ $m\geq 2$ or $s>1$ $A\succeq 0$, (4.1b) an equality, (4.1b) an inequality, (4.1b) an inequality, $\mathbb{P}={\rm St}(k,n)_{D+}$, $\mathbb{P}={\rm St}(k,n)_{D+}$. $\mathbb{P}={\rm St}(k,n)$. or $\mathbb{P}={\rm St}(k,n)$. NEPv $m=s=1$ $m\geq 2$ or $s>1$ (7.1b) an inequality, (7.1b) an equality, (7.1b) an inequality, $\mathbb{P}={\rm St}(k,n)_{D+}$. $\mathbb{P}={\rm St}(k,n)$. $A\succeq 0$, $\mathbb{P}={\rm St}(k,n)$. * Integer $m\geq 1$, scalar $s\geq 1$ and $A$ is symmetric.
@@
and thus $H(P)P-\mathscr{H}(P)\equiv P\big{[}m(D^{\operatorname{T}}P)^{m}\big{]}$ for $P\in{\rm St}(k,n)$. Then we have
$\operatorname{tr}(P^{\operatorname{T}}H(P)P)$ $=2m\,\operatorname{tr}((P^{\operatorname{T}}D)^{m})\quad\mbox{for$P\in{\rm St}(k,n)$};$ (7.7a)
@@
where $\mathbb{P}={\rm St}(k,n)_{D+}$. They, as argued in Remark 7.1, imply that (7.1) holds with $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=2(m-1)$, and thus $f(P)=\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ is an atomic function for NEPv. Furthermore, any solution $P$ to NEPv (6.3) with $H(P)$ in (7.6) such that $(P^{\operatorname{T}}D)^{m}$ is symmetric is a solution to the KKT condition (2.3) and vice versa.
@@
Equation (7.7a) can be straightforwardly verified. Now we prove (7.7b). Inequality (7.7b) for $m=1$ in fact holds for all $P\in{\rm St}(k,n)$. To see this, for $m=1$ and $P\in{\rm St}(k,n)$, since $\widetilde{P}^{\operatorname{T}}D\succeq 0$,
@@
by Lemma B.9. In general for $m>1$, suppose both $P^{\operatorname{T}}D\succeq 0$ and $\widetilde{P}^{\operatorname{T}}D\succeq 0$. Then
@@
Finally (7.7) implies (7.1), as argued in Remark 7.1. ∎
For any $s>1$, $[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ is well-defined for any $P\in\mathbb{R}^{n\times k}$ such that $\operatorname{tr}((P^{\operatorname{T}}D)^{m})\geq 0$. In particular, $[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ is well-defined for $P\in{\rm St}(k,n)_{D+}$. Combining Theorems 7.2 and 7.4, we obtain the following corollary.
@@
Let $D\in\mathbb{R}^{n\times k}$, integer $m\geq 1$, $s>1$, $g(P)=[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$, and $\mathbb{P}={\rm St}(k,n)_{D+}$. Let $H(P)$ be as in (7.6) and $H_{g}(P)=s\,[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s-1}H(P)$ for which $H_{g}(P)P-\partial g(P)/\partial P\equiv P\big{[}sm\,[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s-1}\,(D^{\operatorname{T}}P)^{m}\big{]}$. Then
@@
and thus $\mathscr{H}(P)\equiv H(P)P$ for $P\in\mathbb{R}^{n\times k}$.
@@
They, as argued in Remark 7.1, imply that (7.1) holds with $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=2(m-1)$, and $\mathbb{P}={\rm St}(k,n)$, and thus $f(P)=\operatorname{tr}((P^{\operatorname{T}}AP)^{m})$ is an atomic function for NEPv.
@@
They, as argued in Remark 7.1, imply that (7.1) holds with $\underaccent{\bar}{\alpha}=2$ and $\underaccent{\bar}{\beta}=2(sm-1)$, and $\mathbb{P}={\rm St}(k,n)$, and thus $g(P)=[\operatorname{tr}((P^{\operatorname{T}}AP)^{m})]^{s}$ for $s>1$ is an atomic function for NEPv.
@@ 8 Convex Composition
We are interested in solving the same optimization problem on the Stiefel manifold ${\rm St}(k,n)$ as in (5.1) by Algorithm 1 and its accelerating variation in Algorithm 2 with convergence guarantee. In that regard, we stick to the initial setup at the beginning of section 5 up to the paragraph containing (5.3). We then go along a different path – the path of the NEPv approach. To that end, we will have to specify what $H(P)$, a symmetric matrix-valued function, to use for a given objective function $f=\phi\circ T$ in (5.1), assuming that a symmetric matrix-valued function has already been constructed for each component of $T(P)$.
@@ 8.1 All $P_{i}$ are the entire $P$
$\operatorname{tr}(P^{\operatorname{T}}H_{i}(P)P)$ $=\underaccent{\bar}{\gamma}_{i}\,f_{i}(P)\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (8.5a)
and given $\widehat{P}\in{\rm St}(k,n)$ and $P\in\mathbb{P}$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
Given $P\in\mathbb{P}\subseteq{\rm St}(k,n)$ and $\widehat{P}\in{\rm St}(k,n)$, suppose that (6.7) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\geq\operatorname{tr}(P^{\operatorname{T}}H(P)P)+\eta$. Write
@@
for which $H(P)P-\mathscr{H}(P)\equiv P\big{[}\mathscr{D}(P)^{\operatorname{T}}P\big{]}$ for $P\in{\rm St}(k,n)$, where, as in (5.14),
@@
Any solution $P$ to NEPv (6.3) with $H(P)$ in (8.6) such that $[\mathscr{D}(P)]^{\operatorname{T}}P$ is symmetric is a solution to the KKT condition (2.3) and vice versa. It can be seen that (8.5b) is an equality for $1\leq i\leq\ell$ and hence it does not need to require $\phi_{i}\geq 0$ for $1\leq i\leq\ell$. In addition to this, instead of treating each $H_{\ell+j}(P)$ separately, we can treat $H_{\ell+j}(P)$ for $1\leq j\leq t$ collectively all at once through $\mathscr{D}(P)$, as we did in (5.2), making all $\phi_{\ell+j}\geq 0$ unnecessary as well. A much more improved version of Theorem 8.1 is stated as Theorem 8.2 below, according to which, the best $Q_{i}$ at Line 4 of Algorithm 1 when applied to $\phi\circ T_{1a}$ is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}\mathscr{D}(P^{(i)})$. A special case of $T_{1a}$ is: $t=0$, $k=1$, $P_{i}=\boldsymbol{p}$ (a unit vector) for $1\leq i\leq\ell$, which gives the main problem of [5] (in the paper, $\phi(\boldsymbol{x})=\sum_{i=1}^{\ell}\psi_{i}(x_{i})$ for $\boldsymbol{x}=[x_{1},x_{2},\ldots,x_{\ell}]^{\operatorname{T}}$ with each $\psi_{i}$ being a convex function of a single-variable).
@@
Consider $f=\phi\circ T_{1a}$, and let $\mathscr{D}(P)$ be as in (5.14) and $H(P)$ as in (8.6). Given $\widehat{P}\in{\rm St}(k,n),\,P\in{\rm St}(k,n)$, let $\widetilde{P}=\widehat{P}Q$ where $Q$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{D}(P)$. If
@@
$T_{2a}\,:\,P\in{\rm St}(k,n)\to T_{2a}(P):=\begin{bmatrix}\|P^{\operatorname{T}}A_{1}P\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P^{\operatorname{T}}A_{\ell}P\|_{\operatorname{F}}^{2}\\ \|P^{\operatorname{T}}D_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P^{\operatorname{T}}D_{t}\|_{\operatorname{F}}^{2}\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (8.8)
@@
for which $\mathscr{H}(P)\equiv H(P)P$ for $P\in\mathbb{R}^{n\times k}$. If all $A_{i}\succeq 0$, then, by Theorem 7.5, the consistency conditions in (8.5) are satisfied with $\mathbb{P}={\rm St}(k,n)$, $Q=I_{k}$, $\underaccent{\bar}{\alpha}=2$, $\underaccent{\bar}{\beta}_{i}=2$ for $1\leq i\leq\ell$ and $\underaccent{\bar}{\beta}_{\ell+j}=0$ for $1\leq j\leq t$. Note, for the example, (8.5b) for $\ell+1\leq i\leq\ell+t$ are equalities and hence Theorem 8.1 requires $\phi_{i}\geq 0$ for $1\leq i\leq\ell$ only. Optimization problem (5.1) with $T=T_{2a}$ for $t=0$ and $\phi(\boldsymbol{x})=\sum_{i=1}^{\ell}x_{i}$ gives the key optimization problem in the uniform multidimensional scaling (UMDS) method [79].
Comparing the conclusion here and that of Example 5.2 (for all $P_{i}=P$), we don’t require $\phi_{\ell+j}\geq 0$ for $1\leq j\leq t$ here, everything else being equal. In particular, both require $A_{i}\succeq 0$ for all $i$. Next, we explain how to make the NEPv approach work on $f=\phi\circ T_{2a}$ even if some $A_{i}\not\succeq 0$, yielding yet another example for which the NEPv approach works but the NPDo approach may not. Let $\delta_{i}\in\mathbb{R}$ such that $\widehat{A}_{i}=A_{i}-\delta_{i}I\succeq 0$ for $1\leq i\leq\ell$. This is always possible by letting $\delta_{i}$ be some lower bound of the eigenvalues of $A_{i}$, and numerically, $\delta_{i}$ can be estimated cheaply [81]. Notice that
$\operatorname{tr}((P^{\operatorname{T}}A_{i}P)^{2})=\operatorname{tr}((P^{\operatorname{T}}\widehat{A}_{i}P)^{2})+2\delta_{i}\operatorname{tr}(P^{\operatorname{T}}\widehat{A}_{i}P)+k\delta_{i}^{2}.$
@@
$\widehat{T}_{2a}\,:\,P\in{\rm St}(k,n)\to\widehat{T}_{2a}(P):=\begin{bmatrix}\operatorname{tr}((P^{\operatorname{T}}\widehat{A}_{1}P)^{2})\\ \vdots\\ \operatorname{tr}((P^{\operatorname{T}}\widehat{A}_{\ell}P)^{2})\\ \operatorname{tr}(P^{\operatorname{T}}\widehat{A}_{1}P)\\ \vdots\\ \operatorname{tr}(P^{\operatorname{T}}\widehat{A}_{\ell}P)\\ \|P^{\operatorname{T}}D_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P^{\operatorname{T}}D_{t}\|_{\operatorname{F}}^{2}\\ \end{bmatrix}\in\mathbb{R}^{2\ell+t},$ (8.10)
@@
$T_{3a}\,:\,P\in{\rm St}(k,n)\to T_{3a}(P):=\begin{bmatrix}\operatorname{tr}((P^{\operatorname{T}}A_{1}P)^{m_{1}})\\ \vdots\\ \operatorname{tr}((P^{\operatorname{T}}A_{\ell}P)^{m_{\ell}})\\ \operatorname{tr}((P^{\operatorname{T}}D)^{m_{\ell+1}})\end{bmatrix}\in\mathbb{R}^{\ell+1},$ (8.11)
@@
for which $H(P)P-\mathscr{H}(P)\equiv P\big{[}\phi_{\ell+1}(T_{3a}(P))\,m_{\ell+1}\,(D^{\operatorname{T}}P)^{m_{\ell+1}}\big{]}$ for $P\in{\rm St}(k,n)$. Any solution $P$ to NEPv (6.3) with $H(P)$ in (8.12) such that $(D^{\operatorname{T}}P)^{m_{\ell+1}}$ is symmetric is a solution to the KKT condition (2.3). Suppose all $A_{i}\succeq 0$ and let $\mathbb{P}={\rm St}(k,n)_{D+}$. Then the consistency conditions in (8.5) are satisfied with $\underaccent{\bar}{\alpha}=2$, $\underaccent{\bar}{\beta}_{i}=2(m_{i}-1)$ for $11\leq i\leq\ell+$ 1, by Theorems 7.4 and 7.5. In applying Theorem 8.1 with $T_{0}=T_{3a}$ we need $\phi_{i}\geq 0$ for $11\leq i\leq\ell+$ 1 because now we are not sure if any of the inequalities in (8.5b) is an equality. Lastly, the best $Q_{i}$ at Line 4 of Algorithm 1 when applied to $\phi\circ T_{3a}$ is an orthogonal polar factor of $[\widehat{P}^{(i)}]^{\operatorname{T}}D$ and vice versa.